Don is 5 years older than Peter is. In 5 years, Don will be twice as old as Peter is now. How old is Peter now?
Answer by ikleyn(52834)

.
D = P + 5,           (1)     ("Don is 5 years older than Peter is")
D + 5 = 2*P          (2)     ("In 5 years, Don will be twice as old as Peter is now.")


From (1), substitute D = P+5 into equation (2). You will get


(P+5) + 5 = 2*P  ====>  P = 10.


Answer.  Peter is 10 years old.  Don is 10+5 = 15 years old.